# Unusual critical points between atomic insulating phases

### v1

@@ III.2 Bosonic Atomic Insulator on the Breathing Kagome Lattice
where we have defined translation $T_{1}$ ($T_{2}$ acts trivially on $\Psi$), reflection $\mathcal{R}_{x}$ taking $x\rightarrow-x$, and $C_{3}$ rotation in addition to time reversal $\mathcal{T}$. The above symmetries rule out any
@@ III.2.1 A Review of the Critical Theory
The quantum field theory in Eq. (24) has been studied extensively, and we will briefly review the relevant details here. We first remark that without $\delta{\mathcal{L}}_{E}$, it is believed that Eq. (24) flows to an interacting CFT fixed point Appelquist et al. (1988); Chester and Pufu (2016); Li (2022a); Albayrak et al. (2022); He et al. (2022); Li (2022b) for $N_{f}\gtrsim 4$. Therefore, to analyze the QED3 critical point, one must classify the symmetry-allowed relevant operators contained in $\delta\mathcal{L}_{E}$.
As the Dirac fermion $\psi$ is not a local observable, the most important class of local operators are monopole operators $\mathcal{M}_{q}$, which insert an integer $q$ units of $U(1)$ gauge flux at a spacetime point and describe topologically nontrivial configurations of the gauge field Borokhov et al. (2002a, b). The monopole operators are charged under the $U(1)_{top}$ conserved current
@@
All other local operators, such as fermion billinears, can be constructed as composites of the monopole operators (for a more detailed discussion, one can refer to the Appendix of Zhang et al. (2025a)). Therefore, the monopole operators can be viewed as the core building blocks of the theory. Furthermore, from a $1/N_{f}$ expansion, the scaling dimension of the fundamental monopoles $\phi$ is given by Pufu (2014); Dyer et al. (2013)
@@
which is relevant at the QED3 fixed point. Barring fermion billinears, we will assume all other operators (including four fermion terms and higher charge monopoles) in the theory to be irrelevant. These assumptions are within those commonly made for studies of QED3 theories arising from the lattice He et al. (2022); Zhang et al. (2025a, b). Therefore, the fundamental monopoles are the most important operators in determining the ultimate fate of the quantum critical point. Using the methods developed in Song et al. (2020, 2019), we will analyze the quantum numbers of monopole operators.
Before proceeding, we note that Eq. (24) will also generically contain velocity anisotropy terms. Refs. Franz et al. (2002); Vafek et al. (2002); Hermele et al. (2005, 2007) have shown a class of these, specified by a valley anisotropy, are irrelevant. Other velocity anisotropies include intravalley contributions. Using a large $N_{f}$ expansion, we explicitly show that all velocity anisotropies are irrelevant at the QED3 fixed point in Appendix C.
@@ III.2.2 Monopole Quantum Numbers
To find the Berry phase associated with $\mathcal{T}_{IR}$ and $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$, we follow the analytic argument from Song et al. (2020). For simplicity, let us assume the UV lattice system has a full $SO(3)_{f}$ pseudospin symmetry. First, we introduce a Dirac quantum spin Hall mass $\overline{\psi}\sigma^{3}\psi$, which splits the zero mode degeneracy and leaves only the monopole $\Phi\sim\phi_{4}+i\phi_{5}$ that transforms as $S_{z}$ to be gapless. As the spin Hall mass preserves the IR time reversal, the time reversal Berry phase for the original system can be calculated in the presence of this spin Hall mass. Then, the nontriviality of the spin Hall state $\mathbb{Z}_{2}$ topological insulator Kane and Mele (2005) is manifested through the monopole acquiring a minus sign under the Kramers time reversal, $\Phi\rightarrow-\Phi^{\dagger}$. This can be seen by examining the edge theory of the spin Hall insulating state and identifying the monopole with a particular tunneling operator on the edge. The same arguments apply to $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$, leading to
@@
To arrive at the transformations, we first note that the translations $T_{1,2}$ should act with trivial $U(1)_{top}$ Berry phase. This can be shown algebraically using that the critical point has an emergent $C_{6}$ about each hexagonal plaquette. From the relation $T_{1}C_{6}T_{2}=T_{2}C_{6}$ and the fact that $\phi_{4,5,6}$ transform with only a phase under $T_{1,2}$ and $C_{6}$, one can see that the Berry phase for $T_{1}$ must vanish. The Berry phase for $T_{2}$ must also vanish by the relation $C_{6}T_{1}=T_{2}C_{6}$. To find the Berry phase for $C_{3}$, we can use numerical methods Song et al. (2019), putting the system on a torus and threading $2\pi$ flux uniformly. Filling the appropriate zero modes (which will have a finite size gap), we calculate the quantum numbers of the resulting many body state to confirm the valley singlet monopoles have zero $C_{3}$ angular momentum, as well as zero lattice momentum.
@@ III.2.3 The bosonic critical point
The next most relevant operators are the fermion billinear operators, which we have tuned to zero to reach the critical point. As discussed earlier, other singlet operators, which include excited charge $1$ monopole operators, higher charge monopoles, and operators which transform as symmetric two-index tensors of $SO(6)$, are likely irrelevant by analytic and numeric computations, or they must be assumed irrelevant if $N_{f}=4$ QED3 is found to be stable in any lattice gauge model (more detailed discussion in Zhang et al. (2025a)). Therefore, we assert that the critical point between the two trivial insulating phases on the breathing kagome lattice is the QED3 CFT. This CFT is stable and unfine-tuned as there is only a single relevant perturbation (the fermion mass) tuning the transition.
The insulators on each side of the transition can be viewed as QED3 perturbed by an $SU(4)$ adjoint mass, which breaks the IR global symmetry to $(SO(4)\times SO(2))/\mathbb{Z}_{2}$. However, note that the appearance of a trivial single-charge monopole is forbidden by the UV symmetries. While it is possible that the adjoint mass can lead to monopole proliferation which precipitates spontaneous symmetry breaking Song et al. (2019), we assume instead that the adjoint mass drives proliferation of higher charge, UV singlet monopoles, which further breaks the IR global symmetry to the UV symmetry of the lattice insulator. These higher charge monopoles, which drive confinement, are in the class of UV singlet operators that are irrelevant at the critical point, but become relevant once we deform away from criticality. The schematic scaling of the fermion (parton) gap, $m$, and the photon gap $\Delta_{ph}$ away from the critical point is shown in Fig. 8. In particular, we note that as the photon is gapped by higher charge, dangerously irrelevant monopoles, $\Delta_{ph}$ will be parametrically smaller than the fermion gap $m$.
@@ IV Anomaly Matching Conditions
In this section, we confirm the quantum numbers of the monopoles in the three models we have analyzed is indeed consistent with what is expected from the lattice Lieb-Shultz-Mattis-Oshikawa-Hastings (LSMOH) constraints Lieb et al. (1961); Oshikawa (2000); Hastings (2004); Cheng et al. (2016); Po et al. (2017b); Huang et al. (2017); Cho et al. (2017); Jian et al. (2018); Thorngren and Else (2018). Applying the tools developed in Zou et al. (2021); Ye et al. (2022), this illustrates that the critical theories discussed and the symmetry transformations of the monopoles, including for the unstable multicritical points on the bipartite lattices, are consistent with what can be emergent from a local bosonic lattice Hamiltonian like the ones we described, providing a nontrivial check on our results.
@@
All the models we consider are particularly simple, in that there is no lattice LSMOH anomaly. Therefore, all of the IR anomalies should pullback to zero and the LHS of Eq. (43) is trivial. However, as the RHS of Eq. (43) describes the nontrivial IR anomaly, the matching condition still enforces nontrivial relations for the symmetry embedding $\varphi$. To explicitly show this, we first write down the IR anomaly for $N_{f}=4$ QED3 by coupling $SO(6)$ and $U(1)_{top}=SO(2)$ gauge fields, considering it as a Stiefel liquid Zou et al. (2021). The resulting nonlinear sigma model has the same symmetries and anomalies as QED3.
@@ V Outlook
Note that our methods can be straightforwardly extended to higher dimensions Lin and Hughes (2018); Luo and Wu (2023), such as for the pyrochlore Ezawa (2018) or cubic lattice obstructed atomic insulator. However, due to the tendency toward deconfinement in higher dimensions, we expect a conformal critical point to be less common. As QED4 coupled to fermions is free in the IR and behaves like a pure gauge theory, one may need to employ parton constructions that involve a nonabelian gauge group. In the setting of more general gauge groups, such as $SU(N_{c})$, there is a conformal window of $N_{f}$ for which there exists a conformal fixed point (the Banks-Zak fixed point Caswell (1974); Banks and Zaks (1982)). For $SU(2)$ fundamental fermions, we must have approximately $8\lesssim N_{f}\lesssim 11$, which is harder to realize in lattice models unless there are multiple degenerate bands. However, these general theories could be an interesting avenue of exploration as $(3+1)d$ gauge theories realize a large class of “unnecessary” quantum critical points Bi and Senthil (2019), which may emerge as low energy descriptions in both fermionic and bosonic lattice insulators.
@@
Lastly, as the models explored in this paper are relatively simple, there is hope that the unusual critical behavior, at least in the non-fined tune case of the breathing kagome lattice, can be probed in both numerical simulations and laboratory experiments.
@@ A.1 The Fermionic BBH Model
The BBH model Benalcazar et al. (2017b, a) consists of four orbitals describing spinless fermions on a square lattice. There are dimerized hopping amplitudes corresponding to intra- and inter-site coupling. Furthermore, $\pi$ flux is threaded through each plaquette as shown in Fig. 5. The hopping Hamiltonian is given by,
@@
Under these symmetries, the only allowed mass term is of the form $(\mu^{2}+\mu^{3})\tau^{3}$, so the Dirac theory is indeed a quantum critical point between two different trivial insulator phases. This theory has been extensively studied due to its emergence in graphene, and short-range interactions are irrelevant at the critical point. Furthermore, transitions in fermionic obstructed atomic insulators arising from the annihilation and reappearance of Dirac cones have been analyzed in Radha and Lambrecht (2021).
@@ A.2 The Bosonic BBH model
We will also maintain full $SU(2)$ pseudospin symmetry throughout, in addition to $U(1)_{b}$ number conservation symmetry of the hardcore bosons, which rotates $d_{1,2}\rightarrow e^{i\sigma^{3}}d_{1,2}$. We observe that the time reversal symmetry is what constrains both fermionic partons to simultaneously go critical. Note that within the decomposition $b=d_{1}^{\dagger}d_{2}$, there is an inherent $SU(2)_{g}$ gauge symmetry $(d_{1}^{\dagger},d_{2})\rightarrow U(d_{1}^{\dagger},d_{2})$. This is the same framework used in parton constructions of spin liquid states Wen (2002). The mean field Hamiltonian $H_{MF}$ actually preserves the full $SU(2)_{g}$ gauge symmetry inherent in the parton description. One way to see this is to rewrite
@@ A.3 Monopole quantum numbers
Now to fully constrain the monopole transformations, especially the Berry phases of the UV lattice symmetry transformations, we will again borrow from Song et al. (2020). Note that due to the particle-hole symmetry inherent in the bipartite square lattice, the hopping Hamiltonian can be adiabatically tuned to a point with an enhanced $SU(2)_{g}$ gauge symmetry, as in the case of taking the diagonal hopping terms $\lambda_{2}\rightarrow 0$. A particle-hole symmetry is then restored at this point. For example, for a site $i_{a}$ on the $a$ sublattice, there is a symmetry
@@
For the symmetries of charge conjugation and reflection, we observe that the operator $\mathcal{C}\mathcal{R}_{x}$ can be viewed as a combination of $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$ and a Lorentz rotation that does not affect the monopoles, which are Lorentz scalars. The overall phase of $\mathcal{C}\mathcal{R}_{x}$ can be fixed by noting that $(\mathcal{C}\mathcal{R}_{x})^{2}=1$ on the fermions. From the arguments in Song et al. (2020), $(\mathcal{C}\mathcal{R}_{x})^{2}=1$ is another manifestation of the nontriviality of the spin Hall insulator, in which the spin triplet monopoles must be odd under $\mathcal{CR}_{x}$. Therefore, the (spin triplet) monopoles in this system must also transform the same way, which uniquely determines the combined action $\mathcal{C}\mathcal{R}_{x}$. Note that one of $\mathcal{C}$ and $\mathcal{R}_{x}$ can be defined with an arbitrary overall $U(1)_{top}$ phase, as this amounts to shifting all the monopoles by an overall phase. However, their relative actions are physically relevant and determined by the arguments above. The symmetry transformations are gathered in Table 2.
@@ A.3.1 Multicriticality
There are also other allowed monopole-antimonopole terms such as $\phi_{1}^{\dagger}\phi_{1}$, and these terms have scaling dimension $\sim 2.38$ from large $N_{f}$. However, as mentioned earlier, we will assume that they are irrelevant as is done in numerical calculations, due to the fact that any lattice realization of QED3 will contain singlets of this form He et al. (2022); Zhang et al. (2025a, b).
@@
for $w_{4}[A^{SO(5)}]\in H^{4}(X_{4},\mathbb{Z}_{2})$, corresponding to a discrete $\theta$-angle for the background $SO(5)$ gauge field $A^{SO(5)}$. Proliferating the second trivial monopole will restrict our symmetry to an $SO(4)$ subgroup of $SO(5)$. Decomposing $SO(4)=(SU(2)_{L}\times SU(2)_{R})/\mathbb{Z}_{2}$, the anomaly becomes Wang et al. (2017)
@@
where $CS_{SU(2),X_{4}}[A]=\frac{1}{4\pi}\int_{X_{4}}\operatorname{tr}_{fund}F\wedge F$ is the extension of the Chern-Simons term into $X_{4}$. This anomaly corresponds to opposite discrete $\theta$-angles for $SU(2)_{L}$ and $SU(2)_{R}\subset SO(4)$, and with time reversal, the $\theta$ angles do not flow. Most importantly, the resulting theory cannot be a trivial, symmetric gapped phase. There is a possibility that in the presence of two singlet monopoles, the critical point flows to a conformal fixed point distinct from QED3. Note that despite the ’t Hooft anomaly, the state enjoys symmetry-enforced gaplessness Wang and Senthil (2014, 2016); Sodemann et al. (2017); Wang et al. (2017); Cordova and Ohmori (2020); Córdova and Ohmori (2020), in which the anomaly precludes even a gapped symmetric topologically ordered state. The argument is similar to the one in Dumitrescu et al. (2024). Because $SU(2)$ is simply-connected, it cannot have a non-trivial action on a $(2+1)d$ TQFT; that is, $SU(2)$ cannot carry any non-trivial anomaly. However, the $\theta$ terms imply a non-trivial anomaly for an $SU(2)$ subgroup of $SO(4)$. Instead, a topologically ordered state must break the $SO(4)$ or time reversal (in which case the $\theta$ can flow to be trivial).
@@
Therefore, as in the case posited for a single monopole Zhang et al. (2025a, b), it is more plausible that the final fate of the theory could be some ordered state that spontaneously breaks the remaining $SO(4)$ symmetry. In either case, the highly multicritical nature of the critical point leads us to conclude the transition we have constructed is likely first-order.
@@
We observe that in the above picture, the two insulating phases of the BBH model near the critical point correspond to turning on a suitable $SU(4)$ adjoint mass for $\psi$ while not proliferating any of the two singlet monopoles. This gaps out the fermion matter fields, leaving a pure gauge theory in the IR. Upon turning on the mass, the degeneracy of the six monopoles will be lifted, and there will be one unique gapless monopole in the IR, specifically $\phi_{1}\pm i\frac{\phi_{2}-\phi_{3}}{\sqrt{2}}$. This monopole will spontaneously proliferate as it is also a singlet under the UV symmetries, leading to confinement of the resulting gauge theory. The resulting state (in addition to the discrete symmetries of $\mathcal{C}_{IR}$, $\mathcal{T}_{IR}$, and $\mathcal{R}_{IR}$) will have an $SO(4)\times SO(2)/\mathbb{Z}_{2}$ symmetry, where $SO(2)$ acts as a combination of $U(1)_{top}$ and rotating among the singlet monopoles while $SO(4)$ rotates the other four and $\mathbb{Z}_{2}$ acts as before. A simple ordered state is allowed, and consequently, there is no anomaly even in the IR. However, we still expect the IR symmetry to completely break the global symmetry to the symmetry of the UV square lattice space group. This is because in the presence of the adjoint mass, the UV singlet terms allowed in $\delta{\mathcal{L}}_{E}$ that were irrelevant at the critical point (such as higher strength monopoles) will become relevant and will completely break the IR continuous vector symmetries into discrete subgroups corresponding to the UV lattice symmetries. We remark that the activation of an adjoint mass in QED3 is also a possible mechanism by which conventional Landau symmetry-breaking orders can originate from the a Dirac spin liquid phase or critical point, also described by QED3 Song et al. (2019). However, our case is distinct in that the adjoint mass is not a mechanism for symmetry breaking as it is already a singlet under all UV symmetries.

### v2

@@ III.2 Bosonic Atomic Insulator on the Breathing Kagome Lattice
where we have defined translation $T_{1}$ ($T_{2}$ acts trivially on $\Psi$), reflection $\mathcal{R}_{x}$ taking $x\rightarrow-x$, and $C_{3}$ rotation in addition to time reversal $\mathcal{T}$.
@@ III.2.1 A Review of the Critical Theory
The quantum field theory in Eq. (24) has been studied extensively, and we will briefly review the relevant details here. We first remark that without $\delta{\mathcal{L}}_{E}$, it is believed that Eq. (24) flows to an interacting CFT fixed point [3, 16, 38, 2, 29, 39] for $N_{f}\gtrsim 4$. Therefore, to analyze the QED3 critical point, one must classify the symmetry-allowed relevant operators contained in $\delta\mathcal{L}_{E}$.
As the Dirac fermion $\psi$ is not a local observable, the most important class of local operators are monopole operators $\mathcal{M}_{q}$, which insert an integer $q$ units of $U(1)$ gauge flux at a spacetime point and describe topologically nontrivial configurations of the gauge field [11, 10]. The monopole operators are charged under the $U(1)_{top}$ conserved current
@@
All other local operators, such as fermion billinears, can be constructed as composites of the monopole operators (for a more detailed discussion, one can refer to the Appendix of [74]). Therefore, the monopole operators can be viewed as the core building blocks of the theory. Furthermore, from a $1/N_{f}$ expansion, the scaling dimension of the fundamental monopoles $\phi$ is given by [49, 22]
@@
which is relevant at the QED3 fixed point. Barring fermion billinears, we will assume all other operators (including four fermion terms and higher charge monopoles) in the theory to be irrelevant. These assumptions are within those commonly made for studies of QED3 theories arising from the lattice [29, 74, 75]. Therefore, the fundamental monopoles are the most important operators in determining the ultimate fate of the quantum critical point. Using the methods developed in [59, 60], we will analyze the quantum numbers of monopole operators.
Before proceeding, we note that Eq. (24) will also generically contain velocity anisotropy terms. Refs. [24, 63, 32, 33] have shown a class of these, specified by a valley anisotropy, are irrelevant. Other velocity anisotropies include intravalley contributions. Using a large $N_{f}$ expansion, we explicitly show that all velocity anisotropies are irrelevant at the QED3 fixed point in Appendix C.
@@ III.2.2 Monopole Quantum Numbers
To find the Berry phase associated with $\mathcal{T}_{IR}$ and $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$, we follow the analytic argument from [59]. For simplicity, let us assume the UV lattice system has a full $SO(3)_{f}$ pseudospin symmetry. First, we introduce a Dirac quantum spin Hall mass $\overline{\psi}\sigma^{3}\psi$, which splits the zero mode degeneracy and leaves only the monopole $\Phi\sim\phi_{4}+i\phi_{5}$ that transforms as $S_{z}$ to be gapless. As the spin Hall mass preserves the IR time reversal, the time reversal Berry phase for the original system can be calculated in the presence of this spin Hall mass. Then, the nontriviality of the spin Hall state $\mathbb{Z}_{2}$ topological insulator [36] is manifested through the monopole acquiring a minus sign under the Kramers time reversal, $\Phi\rightarrow-\Phi^{\dagger}$. This can be seen by examining the edge theory of the spin Hall insulating state and identifying the monopole with a particular tunneling operator on the edge. The same arguments apply to $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$, leading to
@@
To arrive at the transformations, we first note that the translations $T_{1,2}$ should act with trivial $U(1)_{top}$ Berry phase. This can be shown algebraically using that the critical point has an emergent $C_{6}$ about each hexagonal plaquette. From the relation $T_{1}C_{6}T_{2}=T_{2}C_{6}$ and the fact that $\phi_{4,5,6}$ transform with only a phase under $T_{1,2}$ and $C_{6}$, one can see that the Berry phase for $T_{1}$ must vanish. The Berry phase for $T_{2}$ must also vanish by the relation $C_{6}T_{1}=T_{2}C_{6}$. To find the Berry phase for $C_{3}$, we can use numerical methods [60], putting the system on a torus and threading $2\pi$ flux uniformly. Filling the appropriate zero modes (which will have a finite size gap), we calculate the quantum numbers of the resulting many body state to confirm the valley singlet monopoles have zero $C_{3}$ angular momentum, as well as zero lattice momentum.
@@ III.2.3 The bosonic critical point
The next most relevant operators are the fermion billinear operators, which we have tuned to zero to reach the critical point. As discussed earlier, other singlet operators, which include excited charge $1$ monopole operators, higher charge monopoles, and operators which transform as symmetric two-index tensors of $SO(6)$, are likely irrelevant by analytic and numeric computations, or they must be assumed irrelevant if $N_{f}=4$ QED3 is found to be stable in any lattice gauge model (more detailed discussion in [74]). Therefore, we find that the critical point between the two trivial insulating phases on the breathing kagome lattice is a natural candidate for realizing the QED3 CFT. This CFT is stable and unfine-tuned as there is only a single relevant perturbation (the fermion mass) tuning the transition.
The insulators on each side of the transition can be viewed as QED3 perturbed by an $SU(4)$ adjoint mass, which breaks the IR global symmetry to $(SO(4)\times SO(2))/\mathbb{Z}_{2}$. However, note that the appearance of a trivial single-charge monopole is forbidden by the UV symmetries. While it is possible that the adjoint mass can lead to monopole proliferation which precipitates spontaneous symmetry breaking [60], we assume instead that the adjoint mass drives proliferation of higher charge, UV singlet monopoles, which further breaks the IR global symmetry to the UV symmetry of the lattice insulator. These higher charge monopoles, which drive confinement, are in the class of UV singlet operators that are irrelevant at the critical point, but become relevant once we deform away from criticality. The schematic scaling of the fermion (parton) gap, $m$, and the photon gap $\Delta_{ph}$ away from the critical point is shown in Fig. 8. In particular, we note that as the photon is gapped by higher charge, dangerously irrelevant monopoles, $\Delta_{ph}$ will be parametrically smaller than the fermion gap $m$.
@@ IV Anomaly Matching Conditions
In this section, we confirm the quantum numbers of the monopoles in the three models we have analyzed is indeed consistent with what is expected from the lattice Lieb-Shultz-Mattis-Oshikawa-Hastings (LSMOH) constraints [40, 45, 28, 15, 47, 34, 17, 35, 62]. Applying the tools developed in [76, 72], this illustrates that the critical theories discussed and the symmetry transformations of the monopoles, including for the unstable multicritical points on the bipartite lattices, are consistent with what can be emergent from a local bosonic lattice Hamiltonian like the ones we described, providing a nontrivial check on our results.
@@
All the models we consider are particularly simple, in that there is no lattice LSMOH anomaly. Therefore, all of the IR anomalies should pullback to zero and the LHS of Eq. (43) is trivial. However, as the RHS of Eq. (43) describes the nontrivial IR anomaly, the matching condition still enforces nontrivial relations for the symmetry embedding $\varphi$. To explicitly show this, we first write down the IR anomaly for $N_{f}=4$ QED3 by coupling $SO(6)$ and $U(1)_{top}=SO(2)$ gauge fields, considering it as a Stiefel liquid [76]. The resulting nonlinear sigma model has the same symmetries and anomalies as QED3.
@@ V Outlook
Note that our methods can be straightforwardly extended to higher dimensions [41, 42], such as for the pyrochlore [23] or cubic lattice obstructed atomic insulator. However, due to the tendency toward deconfinement in higher dimensions, we expect a conformal critical point to be less common. As QED4 coupled to fermions is free in the IR and behaves like a pure gauge theory, one may need to employ parton constructions that involve a nonabelian gauge group. In the setting of more general gauge groups, such as $SU(N_{c})$, there is a conformal window of $N_{f}$ for which there exists a conformal fixed point (the Banks-Zak fixed point [14, 4]). For $SU(2)$ fundamental fermions, we must have approximately $8\lesssim N_{f}\lesssim 11$, which is harder to realize in lattice models unless there are multiple degenerate bands. However, these general theories could be an interesting avenue of exploration as $(3+1)d$ gauge theories realize a large class of “unnecessary” quantum critical points [9], which may emerge as low energy descriptions in both fermionic and bosonic lattice insulators.
@@
Lastly, whether a specific microscopic Hamiltonian realizes the QED3 critical point or is preempted by a weakly first-order transition or intermediate phase is a dynamical question. Our analysis has identified universal properties and the conditions under which QED3 criticality is allowed, and as the models explored in this paper are relatively simple, there is hope that the unusual critical behavior, at least in the non-finetuned case of the breathing kagome lattice, can be probed in the future through both numerical simulations and laboratory experiments.
@@ A.1 The Fermionic BBH Model
The BBH model [6, 5] consists of four orbitals describing spinless fermions on a square lattice. There are dimerized hopping amplitudes corresponding to intra- and inter-site coupling. Furthermore, $\pi$ flux is threaded through each plaquette as shown in Fig. 5. The hopping Hamiltonian is given by,
@@
Under these symmetries, the only allowed mass term is of the form $(\mu^{2}+\mu^{3})\tau^{3}$, so the Dirac theory is indeed a quantum critical point between two different trivial insulator phases. This theory has been extensively studied due to its emergence in graphene, and short-range interactions are irrelevant at the critical point. Furthermore, transitions in fermionic obstructed atomic insulators arising from the annihilation and reappearance of Dirac cones have been analyzed in [50].
@@ A.2 The Bosonic BBH model
We will also maintain full $SU(2)$ pseudospin symmetry throughout, in addition to $U(1)_{b}$ number conservation symmetry of the hardcore bosons, which rotates $d_{1,2}\rightarrow e^{i\sigma^{3}}d_{1,2}$. We observe that the time reversal symmetry is what constrains both fermionic partons to simultaneously go critical. Note that within the decomposition $b=d_{1}^{\dagger}d_{2}$, there is an inherent $SU(2)_{g}$ gauge symmetry $(d_{1}^{\dagger},d_{2})\rightarrow U(d_{1}^{\dagger},d_{2})$. This is the same framework used in parton constructions of spin liquid states [70]. The mean field Hamiltonian $H_{MF}$ actually preserves the full $SU(2)_{g}$ gauge symmetry inherent in the parton description. One way to see this is to rewrite
@@ A.3 Monopole quantum numbers
Now to fully constrain the monopole transformations, especially the Berry phases of the UV lattice symmetry transformations, we will again borrow from [59]. Note that due to the particle-hole symmetry inherent in the bipartite square lattice, the hopping Hamiltonian can be adiabatically tuned to a point with an enhanced $SU(2)_{g}$ gauge symmetry, as in the case of taking the diagonal hopping terms $\lambda_{2}\rightarrow 0$. A particle-hole symmetry is then restored at this point. For example, for a site $i_{a}$ on the $a$ sublattice, there is a symmetry
@@
For the symmetries of charge conjugation and reflection, we observe that the operator $\mathcal{C}\mathcal{R}_{x}$ can be viewed as a combination of $\mathcal{C}_{IR}\mathcal{R}_{x,IR}$ and a Lorentz rotation that does not affect the monopoles, which are Lorentz scalars. The overall phase of $\mathcal{C}\mathcal{R}_{x}$ can be fixed by noting that $(\mathcal{C}\mathcal{R}_{x})^{2}=1$ on the fermions. From the arguments in [59], $(\mathcal{C}\mathcal{R}_{x})^{2}=1$ is another manifestation of the nontriviality of the spin Hall insulator, in which the spin triplet monopoles must be odd under $\mathcal{CR}_{x}$. Therefore, the (spin triplet) monopoles in this system must also transform the same way, which uniquely determines the combined action $\mathcal{C}\mathcal{R}_{x}$. Note that one of $\mathcal{C}$ and $\mathcal{R}_{x}$ can be defined with an arbitrary overall $U(1)_{top}$ phase, as this amounts to shifting all the monopoles by an overall phase. However, their relative actions are physically relevant and determined by the arguments above. The symmetry transformations are gathered in Table 2.
@@ A.3.1 Multicriticality
There are also other allowed monopole-antimonopole terms such as $\phi_{1}^{\dagger}\phi_{1}$, and these terms have scaling dimension $\sim 2.38$ from large $N_{f}$. However, as mentioned earlier, we will assume that they are irrelevant as is done in numerical calculations, due to the fact that any lattice realization of QED3 will contain singlets of this form [29, 74, 75].
@@
for $w_{4}[A^{SO(5)}]\in H^{4}(X_{4},\mathbb{Z}_{2})$, corresponding to a discrete $\theta$-angle for the background $SO(5)$ gauge field $A^{SO(5)}$. Proliferating the second trivial monopole will restrict our symmetry to an $SO(4)$ subgroup of $SO(5)$. Decomposing $SO(4)=(SU(2)_{L}\times SU(2)_{R})/\mathbb{Z}_{2}$, the anomaly becomes [65]
@@
where $CS_{SU(2),X_{4}}[A]=\frac{1}{4\pi}\int_{X_{4}}\operatorname{tr}_{fund}F\wedge F$ is the extension of the Chern-Simons term into $X_{4}$. This anomaly corresponds to opposite discrete $\theta$-angles for $SU(2)_{L}$ and $SU(2)_{R}\subset SO(4)$, and with time reversal, the $\theta$ angles do not flow. Most importantly, the resulting theory cannot be a trivial, symmetric gapped phase. There is a possibility that in the presence of two singlet monopoles, the critical point flows to a conformal fixed point distinct from QED3. Note that despite the ’t Hooft anomaly, the state enjoys symmetry-enforced gaplessness [66, 67, 58, 65, 19, 18], in which the anomaly precludes even a gapped symmetric topologically ordered state. The argument is similar to the one in [21]. Because $SU(2)$ is simply-connected, it cannot have a non-trivial action on a $(2+1)d$ TQFT; that is, $SU(2)$ cannot carry any non-trivial anomaly. However, the $\theta$ terms imply a non-trivial anomaly for an $SU(2)$ subgroup of $SO(4)$. Instead, a topologically ordered state must break the $SO(4)$ or time reversal (in which case the $\theta$ can flow to be trivial).
@@
Therefore, as in the case posited for a single monopole [74, 75], it is more plausible that the final fate of the theory could be some ordered state that spontaneously breaks the remaining $SO(4)$ symmetry. In either case, the highly multicritical nature of the critical point leads us to conclude the transition we have constructed is likely first-order.
@@
We observe that in the above picture, the two insulating phases of the BBH model near the critical point correspond to turning on a suitable $SU(4)$ adjoint mass for $\psi$ while not proliferating any of the two singlet monopoles. This gaps out the fermion matter fields, leaving a pure gauge theory in the IR. Upon turning on the mass, the degeneracy of the six monopoles will be lifted, and there will be one unique gapless monopole in the IR, specifically $\phi_{1}\pm i\frac{\phi_{2}-\phi_{3}}{\sqrt{2}}$. This monopole will spontaneously proliferate as it is also a singlet under the UV symmetries, leading to confinement of the resulting gauge theory. The resulting state (in addition to the discrete symmetries of $\mathcal{C}_{IR}$, $\mathcal{T}_{IR}$, and $\mathcal{R}_{IR}$) will have an $SO(4)\times SO(2)/\mathbb{Z}_{2}$ symmetry, where $SO(2)$ acts as a combination of $U(1)_{top}$ and rotating among the singlet monopoles while $SO(4)$ rotates the other four and $\mathbb{Z}_{2}$ acts as before. A simple ordered state is allowed, and consequently, there is no anomaly even in the IR. However, we still expect the IR symmetry to completely break the global symmetry to the symmetry of the UV square lattice space group. This is because in the presence of the adjoint mass, the UV singlet terms allowed in $\delta{\mathcal{L}}_{E}$ that were irrelevant at the critical point (such as higher strength monopoles) will become relevant and will completely break the IR continuous vector symmetries into discrete subgroups corresponding to the UV lattice symmetries. We remark that the activation of an adjoint mass in QED3 is also a possible mechanism by which conventional Landau symmetry-breaking orders can originate from the a Dirac spin liquid phase or critical point, also described by QED3 [60]. However, our case is distinct in that the adjoint mass is not a mechanism for symmetry breaking as it is already a singlet under all UV symmetries.